# Semi-Supervised Preference Optimization with Limited Feedback

### v2

@@ Abstract
The field of preference optimization has made outstanding contributions to the alignment of language models with human preferences. Despite these advancements, recent methods still rely heavily on substantial paired (labeled) feedback data, leading to substantial resource expenditures. To address these challenges, we study the problem of Semi-Supervised Preference Optimization (SSPO) in which the idea is to learn from both a small number of pairwise preference labels and a large pool of unpaired samples simultaneously. Our key theoretical contribution proves the existence of an optimal reward threshold capable of separating winning and losing responses with high probability, which enables a principled pseudo-labeling of unpaired data. By leveraging these pseudo-labels, SSPO effectively distills latent preferences from large-scale unpaired data, thus maintaining human alignment while drastically reducing acquisition costs. Extensive experiments across datasets validate this remarkable data efficiency; for instance, SSPO trained with Mistral-7B-Instruct on just 1% of UltraFeedback consistently surpasses strong baselines trained on 10% of UltraFeedback. [^1] https://anonymous.4open.science/r/SSPO-F7FE/
@@ 1 Introduction
Preference optimization (PO) is a pivotal method for aligning large language models (LLMs) with human values and expectations, ensuring that models provide useful, safe, and pleasant outputs (calandriello2024humanalignmentlargelanguage). Without proper alignment, LLMs risk generating misleading or harmful content (song2024preferencerankingoptimizationhuman). In PO, preferences reflect human judgments on the desirability or usefulness of responses, encompassing nuanced aspects such as ethical soundness or honesty (christiano2017deep; ouyang2022traininglanguagemodelsfollow; bai2022traininghelpfulharmlessassistant). Building such preference datasets typically involves labor-intensive human annotation processes (cui2024ultrafeedbackboostinglanguagemodels; karthik2024scalablerankedpreferenceoptimization; shi2024saferinstructaligninglanguagemodels), where annotators craft prompts, gather responses from models (achiam2023gpt; dubey2024llama), and rank them.
The primary bottleneck in PO is its profound reliance on data acquisition. This process is costly; it requires expert labor, averaging 5–10 minutes per comparison, and reaching \$10–30 per data point (bai2022traininghelpfulharmlessassistant; casper2023open). To conquer these limitations, recent work has explored synthetic feedback (huang-etal-2023-learning-preference; zhou2024anyprefer; xu2024dr) and automatic preference prediction using LLMs (he-etal-2024-semi-supervised). While promising, these approaches face quality assurance issues due to the lack of verified ground truth.
A potential solution to this costly data acquisition lies in the vast amount of existing domain-specific data, such as question-answering pairs originally intended for supervised fine-tuning (SFT) (agrawal2024modeling). Although rich in expert knowledge, this data lacks the explicit preference labels required for PO. But it often contains valuable implicit preferences, such as coherent thinking patterns and appropriate stylistic tones. To leverage these unlabeled sources, some studies have attempted to generate preference signals using a capable LLM for self-annotation.
However, relying on synthetic data or LLM-based annotations for human alignment has fundamental limitations. This strategy risks creating a feedback loop in which an imperfectly aligned model propagates its own biases and fails to fully capture the complexity of genuine human preferences, thus generating data of questionable reliability (casper2023open). Such synthetic preferences often lack the subtle, context-dependent nuances that are characteristic of true human judgment. Ultimately, acquiring reliable high-quality preference data, especially in specialized domains, still entails prohibitively expensive expert evaluation, creating a significant barrier to developing models that are knowledgeable and aligned with human values in a safe way (miller2024ai).
@@ 2.1 Preference Optimization
Preference optimization aligns language models with human preferences to improve helpfulness and safety. Proximal Policy Optimization (PPO) (schulman2017proximalpolicyoptimizationalgorithms) introduced a stable reinforcement learning algorithm that prevents large policy shifts during training. This method laid the foundation for preference-based fine-tuning of language models. Reinforcement Learning from Human Feedback (RLHF) (ouyang2022traininglanguagemodelsfollow) enhanced alignment by using human rankings to train a reward model, which then guides policy updates. While effective, RLHF’s reliance on reward models can introduce bias and increase training complexity. To simplify this process, Rank Rewarding from Human Feedback (RRHF) (yuan2023rrhf) and Direct Preference Optimization (DPO) (rafailov2024direct) reduced dependence on reward models. RRHF directly incorporates ranking signals, while DPO frames preference learning as a binary classification task, improving stability and efficiency. Odds Ratio Preference Optimization (ORPO) (hong2024orpomonolithicpreferenceoptimization) and Simple Preference Optimization (SimPO) (meng2024simposimplepreferenceoptimization) further improved efficiency by removing the need for a reference model. ORPO models preferences through odds ratios, while SimPO focuses on simplicity and ease of implementation. Kahneman-Tversky Optimization (KTO) (ethayarajh2024ktomodelalignmentprospect) takes inspiration from behavioral economics, incorporating human-like biases (e.g., dislike of loss) into the optimization process to better reflect real-world preferences.
@@ 2.2 Human Alignment with Limited Feedback
Alignment with insufficient feedback is a key challenge in fine-tuning LLMs, where high-quality human preference data is often limited or costly (casper2023open). Ziegler demonstrated the effectiveness of preference modeling for alignment but relied heavily on extensive human annotations. To reduce this dependency, kim2025spreadpreferenceannotationdirect and huang-etal-2023-learning-preference proposed generating synthetic preference data, and zhou2024anyprefer extended this with a generalizable synthesis framework. shi2024saferinstructaligninglanguagemodels and liu2023aligning explored automated preference signals, while Semi-Supervised Reward Modeling (SSRM) (he-etal-2024-semi-supervised) investigates reward modeling using a self-training framework under a semi-supervised setting.
@@ 2.3 Synthetic Preference Generation
Synthetic preference generation reduces dependence on manual labeling but is intrinsically limited by the alignment quality of the generating models. For example, AlpacaFarm (dubois2023alpacafarm) uses LLMs to simulate human preferences, improving cost efficiency but risking the propagation of model biases, which creates a circular constraint on alignment. Meanwhile, off-the-shelf supervised fine-tuning (SFT) data contains valuable implicit preference signals, such as reasoning and stylistic cues, yet is often underutilized. Self-training approaches leverage this data via pseudo-labeling (wang2024self), but typically require iterative annotation cycles that may introduce noise and instability. In comparison, Spread Preference Annotation (SPA) (kim2025spreadpreferenceannotationdirect) addresses scarcity by repeatedly self-annotating with improved preference models to refine alignment progressively, though this process can be computationally intensive and may amplify errors.
@@ 3 Preliminaries
To model our preference classifier, we define $f_{\theta}(x,y,y^{\prime})$ by Bradley-Terry modelization (bradley1952rank; rafailov2024direct; meng2024simposimplepreferenceoptimization):
@@
This formulation captures the expected agreement of the classifier with the given preference label and serves as the foundation for reward-based preference learning. In our study, we set $r_{\theta}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)$ as the default choice, which is the reward function of SimPO (meng2024simposimplepreferenceoptimization). Therefore, the expected risk of the paired data is:
@@
where $\pi_{\theta}(y\mid x)$ is the language model’s policy (likelihood), $\beta$ is a scaling hyperparameter and $\Delta$ is a reward margin guarantee for winning to exceed losing (firth2005bradley; meng2024simposimplepreferenceoptimization).
@@ 4.2 Practical Approach
While Eq. (8) guarantees that an optimal threshold $\delta^{*}$ exists within a certain range, it depends on the unknown means $\mu_{l}$ and $\mu_{w}$. Therefore, we cannot compute $\delta^{*}$ directly in practice. To address this limitation, we estimate the reward densities of winning and losing responses in $D_{L}$ via kernel density estimation (parzen1962estimation; silverman2018density) and solve for the threshold directly:
@@ 5.1 Toy Experiment
To analyze SSPO’s behavior, we design two synthetic experiments: one noise-free and one with label noise. These experiments serve as a sanity check and provide interpretable insights before applying SSPO to real-world preference datasets. Each prompt consists of ten words randomly sampled from the NLTK words (loper2002nltk), with the shortest word designated as preferred over the longest. We fix the unpaired dataset size at 1,000 samples and vary the paired dataset size (10, 50, 100) to simulate data-scarce scenarios. Using GPT-2-small (radford2019language) as a reward model, we compare SSPO against strong baselines under various prior values (0.1, 0.3, 0.5, 0.7, 0.9). To evaluate robustness, the noisy setting swaps the designations of winning and losing for 10%, 30%, and 50% of the paired preference labels.
@@ 5.2 Real-Data Experiments
We conduct training using UltraFeedback (cui2024ultrafeedbackboostinglanguagemodels) for paired preference data. To simulate data-scarce scenarios, we evaluate on two challenging settings: using only 1% ($n_{L}$=611) or 10% ($n_{L}$=6,113) of each original dataset, including validation data. For SSPO, we use 10% of UltraChat-200k (ding2023enhancingchatlanguagemodels) ($n_{U}$=20,786) as the unpaired data to be pseudo-labeled. We utilize off-the-shelf models already fine-tuned on the UltraChat dataset as initial models. We employ Phi-2 (2.7B) (javaheripi2023phi), Mistral-7B-Instruct-v0.2 (jiang2023mistral7b), and Llama3-8B-Instruct (dubey2024llama). to demonstrate the generality and scalability of SSPO.
We also conduct two domain-specific experiments to investigate the alignment with both domain expertise and human preferences. For the medical domain, we set two data sizes according to UltraFeedback; we set 1% ($n_{L}$=1,093) and 10% ($n_{L}$=10,935) of UltraMedical-Preference including validation data, which are composed of pairwise comparisons constructed from real-world medical questions and multiple answers. For SSPO, we use 5% of UltraMedical (zhang2024ultramedical) ($n_{U}$=20,479) as the unpaired data to match the amount of data in accordance with UltraChat. We train Meerkat-7B-v1.0 (kim2024meerkat) and Llama-3-8B-UltraMedical (zhang2024ultramedical) on these datasets. Similarly, for the business domain, we set 1% ($n_{L}$=125) and 10% ($n_{L}$=1,255) of the DSP Business (cheng2023dsp), which provides business-oriented pairwise comparisons. For SSPO, we use whole 17k Business Book ($n_{U}$=17,480) as the unpaired data. We train Mistral-7B-Business and Finance-Llama-8B on these datasets. For more implementation details and hyperparameter settings, please refer to Appendix C.3.
@@
We adopt AlpacaEval2.0 (dubois2025lengthcontrolledalpacaevalsimpleway) and MT-Bench (zheng2023judgingllmasajudgemtbenchchatbot), two widely used benchmarks for assessing models’ conversational and instruction-following abilities across diverse topics and domains. We use GPT-4-Turbo (achiam2023gpt) as the judge model and report both a raw win rate (WR) and a length-controlled win rate (LC), where the LC is the variant of WR explicitly controlling the response-length differences. This correction is designed to address the well-established verbosity bias in LLM-based evaluators, which often exhibit a systematic tendency to favor excessively long outputs irrespective of substantive quality. By conditioning on equal-length responses, LC yields preference estimates that more faithfully reflect semantic adequacy and instruction compliance. Meanwhile, MT-Bench is made up of 80 questions from 8 categories, and we investigate the average score of two turns on a scale of 10 using GPT-4 (achiam2023gpt) as the judge model. MT-Bench provides a complementary perspective by evaluating multi-turn dialog coherence, instruction adherence, and reasoning depth under carefully curated human-written prompts, making it a strong indicator of real-world conversational robustness.
@@ 5.3 Main Results and Ablation Studies
Robustness against Confirmation Bias and Non-stationary Rewards. To address the risks of confirmation bias and reward non-stationarity, we analyze the evolution of reward distributions and the estimated threshold in Figure 3. In the early stages (e.g., step 100), where winning and losing distributions exhibit significant overlap, the adaptive scheduler $\gamma^{\prime}$ prioritizes the supervised loss from $D_{L}$. This acts as a reliable anchor to prevent noise amplification from incorrect pseudo-labels, aligning with the early learning phenomenon where models learn clean patterns before overfitting to noise (arazo2020pseudolabelingconfirmationbiasdeep). As training progresses and a clear margin emerges, SSPO addresses the shifting nature of reward values by estimating the optimal boundary via KDE and stabilizing it with EMA. Drawing on consistency regularization principles (tarvainen2018meanteachersbetterrole), this dynamic thresholding ensures the decision boundary $\hat{\delta}$ consistently tracks the optimal separation without collapsing, enabling the model to effectively leverage unpaired data $D_{U}$ for generalization.
@@ B.2 Gradient Analysis
We define our paired-data risk as Eq. (4). The gradient weight increases when the model incorrectly assigns a higher reward to the losing response. This encourages the model to correct such misrankings by increasing the winning score and decreasing the losing one. This analysis is basically the same as DPO (rafailov2024direct) and SimPO (meng2024simposimplepreferenceoptimization).

### v3

@@ Abstract
The field of preference optimization has made outstanding contributions to the alignment of language models with human preferences. Despite these advancements, recent methods still rely heavily on substantial paired (labeled) feedback data, leading to substantial resource expenditures. To address these challenges, we study the problem of Semi-Supervised Preference Optimization (SSPO) in which the idea is to learn from both a small number of pairwise preference labels and a large pool of unpaired samples simultaneously. Our key theoretical contribution proves the existence of an optimal reward threshold capable of separating winning and losing responses with high probability, which enables a principled pseudo-labeling of unpaired data. By leveraging these pseudo-labels, SSPO effectively distills latent preferences from large-scale unpaired data, thus maintaining human alignment while drastically reducing acquisition costs. Extensive experiments across datasets validate this remarkable data efficiency; for instance, SSPO trained with Mistral-7B-Instruct on just 1% of UltraFeedback consistently surpasses strong baselines trained on 10% of UltraFeedback. [^1] https://github.com/MLAI-Yonsei/SSPO
@@ 1 Introduction
Preference optimization (PO) is a pivotal method for aligning large language models (LLMs) with human values and expectations, ensuring that models provide useful, safe, and pleasant outputs (Calandriello et al., 2024). Without proper alignment, LLMs risk generating misleading or harmful content (Song et al., 2024). In PO, preferences reflect human judgments on the desirability or usefulness of responses, encompassing nuanced aspects such as ethical soundness or honesty (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022). Building such preference datasets typically involves labor-intensive human annotation processes (Cui et al., 2024; Karthik et al., 2024; Shi et al., 2024), where annotators craft prompts, gather responses from models (Achiam et al., 2023; Dubey et al., 2024), and rank them.
The primary bottleneck in PO is its profound reliance on data acquisition. This process is costly; it requires expert labor, averaging 5–10 minutes per comparison, and reaching \$10–30 per data point (Bai et al., 2022; Casper et al., 2023). To conquer these limitations, recent work has explored synthetic feedback (Huang et al., 2023; Zhou et al., 2024; Xu et al., 2024) and automatic preference prediction using LLMs (He et al., 2024). While promising, these approaches face quality assurance issues due to the lack of verified ground truth.
A potential solution to this costly data acquisition lies in the vast amount of existing domain-specific data, such as question-answering pairs originally intended for supervised fine-tuning (SFT) (Agrawal et al., 2024). Although rich in expert knowledge, this data lacks the explicit preference labels required for PO. But it often contains valuable implicit preferences, such as coherent thinking patterns and appropriate stylistic tones. To leverage these unlabeled sources, some studies have attempted to generate preference signals using a capable LLM for self-annotation.
However, relying on synthetic data or LLM-based annotations for human alignment has fundamental limitations. This strategy risks creating a feedback loop in which an imperfectly aligned model propagates its own biases and fails to fully capture the complexity of genuine human preferences, thus generating data of questionable reliability (Casper et al., 2023). Such synthetic preferences often lack the subtle, context-dependent nuances that are characteristic of true human judgment. Ultimately, acquiring reliable high-quality preference data, especially in specialized domains, still entails prohibitively expensive expert evaluation, creating a significant barrier to developing models that are knowledgeable and aligned with human values in a safe way (Miller et al., 2024).
@@ 2.1 Preference Optimization
Preference optimization aligns language models with human preferences to improve helpfulness and safety. Proximal Policy Optimization (PPO) (Schulman et al., 2017) introduced a stable reinforcement learning algorithm that prevents large policy shifts during training. This method laid the foundation for preference-based fine-tuning of language models. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) enhanced alignment by using human rankings to train a reward model, which then guides policy updates. While effective, RLHF’s reliance on reward models can introduce bias and increase training complexity. To simplify this process, Rank Rewarding from Human Feedback (RRHF) (Yuan et al., 2023) and Direct Preference Optimization (DPO) (Rafailov et al., 2024) reduced dependence on reward models. RRHF directly incorporates ranking signals, while DPO frames preference learning as a binary classification task, improving stability and efficiency. Odds Ratio Preference Optimization (ORPO) (Hong et al., 2024) and Simple Preference Optimization (SimPO) (Meng et al., 2024) further improved efficiency by removing the need for a reference model. ORPO models preferences through odds ratios, while SimPO focuses on simplicity and ease of implementation. Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) takes inspiration from behavioral economics, incorporating human-like biases (e.g., dislike of loss) into the optimization process to better reflect real-world preferences.
@@ 2.2 Human Alignment with Limited Feedback
Alignment with insufficient feedback is a key challenge in fine-tuning LLMs, where high-quality human preference data is often limited or costly (Casper et al., 2023). Ziegler et al. (2019) demonstrated the effectiveness of preference modeling for alignment but relied heavily on extensive human annotations. To reduce this dependency, Kim et al. (2025) and Huang et al. (2023) proposed generating synthetic preference data, and Zhou et al. (2024) extended this with a generalizable synthesis framework. Shi et al. (2024) and Liu et al. (2023) explored automated preference signals, while Semi-Supervised Reward Modeling (SSRM) (He et al., 2024) investigates reward modeling using a self-training framework under a semi-supervised setting.
@@ 2.3 Synthetic Preference Generation
Synthetic preference generation reduces dependence on manual labeling but is intrinsically limited by the alignment quality of the generating models. For example, AlpacaFarm (Dubois et al., 2023) uses LLMs to simulate human preferences, improving cost efficiency but risking the propagation of model biases, which creates a circular constraint on alignment. Meanwhile, off-the-shelf supervised fine-tuning (SFT) data contains valuable implicit preference signals, such as reasoning and stylistic cues, yet is often underutilized. Self-training approaches leverage this data via pseudo-labeling (Wang et al., 2024), but typically require iterative annotation cycles that may introduce noise and instability. In comparison, Spread Preference Annotation (SPA) (Kim et al., 2025) addresses scarcity by repeatedly self-annotating with improved preference models to refine alignment progressively, though this process can be computationally intensive and may amplify errors.
@@ 3 Preliminaries
To model our preference classifier, we define $f_{\theta}(x,y,y^{\prime})$ by Bradley-Terry modelization (Bradley and Terry, 1952; Rafailov et al., 2024; Meng et al., 2024):
@@
This formulation captures the expected agreement of the classifier with the given preference label and serves as the foundation for reward-based preference learning. In our study, we set $r_{\theta}(x,y)=\frac{\beta}{|y|}\log\pi_{\theta}(y\mid x)$ as the default choice, which is the reward function of SimPO (Meng et al., 2024). Therefore, the expected risk of the paired data is:
@@
where $\pi_{\theta}(y\mid x)$ is the language model’s policy (likelihood), $\beta$ is a scaling hyperparameter and $\Delta$ is a reward margin guarantee for winning to exceed losing (Firth, 2005; Meng et al., 2024).
@@ 4.2 Practical Approach
While Eq. (8) guarantees that an optimal threshold $\delta^{*}$ exists within a certain range, it depends on the unknown means $\mu_{l}$ and $\mu_{w}$. Therefore, we cannot compute $\delta^{*}$ directly in practice. To address this limitation, we estimate the reward densities of winning and losing responses in $D_{L}$ via kernel density estimation (Parzen, 1962; Silverman, 2018) and solve for the threshold directly:
@@ 5.1 Toy Experiment
To analyze SSPO’s behavior, we design two synthetic experiments: one noise-free and one with label noise. These experiments serve as a sanity check and provide interpretable insights before applying SSPO to real-world preference datasets. Each prompt consists of ten words randomly sampled from the NLTK words (Loper and Bird, 2002), with the shortest word designated as preferred over the longest. We fix the unpaired dataset size at 1,000 samples and vary the paired dataset size (10, 50, 100) to simulate data-scarce scenarios. Using GPT-2-small (Radford et al., 2019) as a reward model, we compare SSPO against strong baselines under various prior values (0.1, 0.3, 0.5, 0.7, 0.9). To evaluate robustness, the noisy setting swaps the designations of winning and losing for 10%, 30%, and 50% of the paired preference labels.
@@ 5.2 Real-Data Experiments
We conduct training using UltraFeedback (Cui et al., 2024) for paired preference data. To simulate data-scarce scenarios, we evaluate on two challenging settings: using only 1% ($n_{L}$=611) or 10% ($n_{L}$=6,113) of each original dataset, including validation data. For SSPO, we use 10% of UltraChat-200k (Ding et al., 2023) ($n_{U}$=20,786) as the unpaired data to be pseudo-labeled. We utilize off-the-shelf models already fine-tuned on the UltraChat dataset as initial models. We employ Phi-2 (2.7B) (Javaheripi et al., 2023), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), and Llama3-8B-Instruct (Dubey et al., 2024). to demonstrate the generality and scalability of SSPO.
We also conduct two domain-specific experiments to investigate the alignment with both domain expertise and human preferences. For the medical domain, we set two data sizes according to UltraFeedback; we set 1% ($n_{L}$=1,093) and 10% ($n_{L}$=10,935) of UltraMedical-Preference including validation data, which are composed of pairwise comparisons constructed from real-world medical questions and multiple answers. For SSPO, we use 5% of UltraMedical (Zhang et al., 2024) ($n_{U}$=20,479) as the unpaired data to match the amount of data in accordance with UltraChat. We train Meerkat-7B-v1.0 (Kim et al., 2024) and Llama-3-8B-UltraMedical (Zhang et al., 2024) on these datasets. Similarly, for the business domain, we set 1% ($n_{L}$=125) and 10% ($n_{L}$=1,255) of the DSP Business (Cheng et al., 2023), which provides business-oriented pairwise comparisons. For SSPO, we use whole 17k Business Book ($n_{U}$=17,480) as the unpaired data. We train Mistral-7B-Business and Finance-Llama-8B on these datasets. For more implementation details and hyperparameter settings, please refer to Appendix C.3.
@@
We adopt AlpacaEval2.0 (Dubois et al., 2025) and MT-Bench (Zheng et al., 2023), two widely used benchmarks for assessing models’ conversational and instruction-following abilities across diverse topics and domains. We use GPT-4-Turbo (Achiam et al., 2023) as the judge model and report both a raw win rate (WR) and a length-controlled win rate (LC), where the LC is the variant of WR explicitly controlling the response-length differences. This correction is designed to address the well-established verbosity bias in LLM-based evaluators, which often exhibit a systematic tendency to favor excessively long outputs irrespective of substantive quality. By conditioning on equal-length responses, LC yields preference estimates that more faithfully reflect semantic adequacy and instruction compliance. Meanwhile, MT-Bench is made up of 80 questions from 8 categories, and we investigate the average score of two turns on a scale of 10 using GPT-4 (Achiam et al., 2023) as the judge model. MT-Bench provides a complementary perspective by evaluating multi-turn dialog coherence, instruction adherence, and reasoning depth under carefully curated human-written prompts, making it a strong indicator of real-world conversational robustness.
@@ 5.3 Main Results and Ablation Studies
Robustness against Confirmation Bias and Non-stationary Rewards. To address the risks of confirmation bias and reward non-stationarity, we analyze the evolution of reward distributions and the estimated threshold in Figure 3. In the early stages (e.g., step 100), where winning and losing distributions exhibit significant overlap, the adaptive scheduler $\gamma^{\prime}$ prioritizes the supervised loss from $D_{L}$. This acts as a reliable anchor to prevent noise amplification from incorrect pseudo-labels, aligning with the early learning phenomenon where models learn clean patterns before overfitting to noise (Arazo et al., 2020). As training progresses and a clear margin emerges, SSPO addresses the shifting nature of reward values by estimating the optimal boundary via KDE and stabilizing it with EMA. Drawing on consistency regularization principles (Tarvainen and Valpola, 2018), this dynamic thresholding ensures the decision boundary $\hat{\delta}$ consistently tracks the optimal separation without collapsing, enabling the model to effectively leverage unpaired data $D_{U}$ for generalization.
@@ 6 Conclusion
Acknowledgement
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00457216).
@@ B.2 Gradient Analysis
We define our paired-data risk as Eq. (4). The gradient weight increases when the model incorrectly assigns a higher reward to the losing response. This encourages the model to correct such misrankings by increasing the winning score and decreasing the losing one. This analysis is basically the same as DPO (Rafailov et al., 2024) and SimPO (Meng et al., 2024).